OpenStudy (anonymous):

Where does the graph of y = 2x2 + 5x − 3 cross the x-axis?

OpenStudy (cwrw238):

it crosses the x-axis when y = 0 so solving 2x^2 + 5x - 3 = 0 for x gives you the values

OpenStudy (cwrw238):

can you solve this?

OpenStudy (anonymous):

me+math= stinky.

OpenStudy (cwrw238):

lol factor the equation (2x - 1 )(x + 3 ) = 0 2x - 1 = 0 or x + 3 = 0 can you solve these 2 equations?

OpenStudy (cwrw238):

if x+ 3 = 0 what is x?

OpenStudy (cwrw238):

no? yes?